# Modeling Tidal Disruptions with Dynamical Tides

### v1

@@ Abstract
Tidal disruption events (TDEs) occur when stars pass close enough to supermassive black holes to be torn apart by tidal forces. Traditionally, these events are studied with computationally intensive hydrodynamical simulations. In this paper, we present a fast, physically motivated two-stage model for TDEs. In the first stage, we model the star’s tidal deformation using linear stellar perturbation theory, treating the star as a collection of driven harmonic oscillators. When the tidal energy exceeds a fraction $\gamma$ of the star’s gravitational binding energy (with $\gamma\sim\mathcal{O}(1)$), we transition to the second stage, where we model the disrupted material as free particles. The parameter $\gamma$ is determined with a one-time calibration to hydrodynamical simulations. This method enables fast computation of the energy distribution $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}E$ and fallback rate $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}T$, while offering physical insight into the disruption process. We apply our model to MESA-generated profiles of middle-age main-sequence stars. Our code is available on GitHub.
@@ 1 Introduction
As famously argued by Frank & Rees (1976) and Magorrian & Tremaine (1999), the TDE rate per galaxy per year is about $10^{-5}$ to $10^{-4}$, which is consistent with current observations (Donley et al. (2002); van Velzen (2018); Sunyaev et al. (2021); Yao et al. (2023)). To date, more than $100$ TDE flares have been observed in the central regions of various galaxies, and this number is expected to increase dramatically in the next few years (Bricman & Gomboc (2019)). Modern observations can now routinely capture both the rising and the decaying parts of the transient light curve (Holoien et al. (2019a); Shappee et al. (2019); van Velzen et al. (2019); Holoien et al. (2019b); Hung et al. (2020)). Once the light curve is observed, we are in principle able to infer the properties of the central BH and of the disrupted stars, e.g. BH mass and spin, stellar mass and age, impact parameter, etc. This information may be used to determine the SMBH mass function and the stellar dynamics near the galactic center (Stone & Metzger (2016); Yao et al. (2023)). In order to extract these properties from the observed data, a great amount of hydrodynamical simulations have been conducted to study the dynamics of TDE (Nolthenius & Katz (1982); Khokhlov et al. (1993a, b); Laguna et al. (1994); Diener et al. (1997); Ramirez-Ruiz & Rosswog (2009); Guillochon & Ramirez-Ruiz (2013); Law-Smith et al. (2017); Golightly et al. (2019b); Goicovic et al. (2019)). Pioneering simulations used polytropic stars models for simplicity, see e.g. Guillochon & Ramirez-Ruiz (2013). More recent simulations have instead studied TDEs of main-sequence stars and included relativistic effects, which presumably play an important role for “deep encounters” (Gafton & Rosswog (2019)), where the star travels very close to the BH horizon.
@@
The challenges in the simulation thus provide a strong motivation to come up with a simplified TDE model that is computationally cheap and capable of capturing the essential physical processes. In Lodato et al. (2009); Kesden (2012a, b); Servin & Kesden (2017); Coughlin & Nixon (2019); Golightly et al. (2019a); Coughlin & Nixon (2022), various approximation methods have been used to make predictions for TDE. For example, the star is assumed to be in hydrostatic equilibrium throughout and follow a nearly parabolic orbit until it reaches the pericenter (alternatively, the tidal radius). Then the star is disrupted instantaneously and each fluid element is modeled as a free particle, which follows its own ballistic orbit in the gravitational field of the black hole. This simplified model gives good qualitative agreement with the simulations.
@@
where $\gamma$ is a calibration factor that can be pre-computed by fitting to hydrodynamical simulations, as it only depends on the stellar structure. The deformed density profile can be evaluated by integrating the tidal response over the star’s trajectory. When Eq. (5) is satisfied, the stellar perturbations are no longer stable, and the star is disrupted. In the second stage, we treat each fluid element inside the star as a free-falling particle, which allows us to evaluate the distribution of the orbital energy of the debris and the mass fallback rate. Even though most of the calculations in this model need to be done numerically, the whole computation only takes around one minute, much cheaper compared with the full hydrodynamical simulation.
@@ 2.2 Tidal energy and disruption criterion
The depth of the potential well is given by the gravitational binding energy of the star $U_{\rm bind}$. We thus expect the star to be disrupted when $E_{Q}\sim|U_{\rm bind}|$. Because we extrapolate our linear oscillations model until the star is disrupted, we allow for a $\mathcal{O}(1)$ correction to the above relation. We thus introduce a calibration factor $\gamma$ and define the tidal disruption criterion as
@@ 2.3 Stage 2: free particles
$\boldsymbol{a}=-\frac{\nabla_{\boldsymbol{x}}p}{\rho}-\nabla_{\boldsymbol{x}}(\Phi_{\star}+\Phi_{\rm BH})$ (36)
@@
Any function $f(E,L_{z})$ of the orbital energy $E$ and angular momentum $L_{z}$ is a solution to Eq. (35) because of the conservation laws for Newtonian orbits.
In practice, we calculate the Lagrangian displacement at the disruption time $t_{\rm TDE}$, which also allows us to obtain the density profile at that instant. After that, we sample the stellar fluid elements to determine the orbital energy distribution $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}E$. In Newtonian gravity, we can further use the relation
@@
to calculate the mass fallback rate $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}T$, which is proportional to the light curve that can be observed (Coughlin & Nixon (2022)).
@@ 3.1 MESA stars on parabolic orbits
The model described in Section 2 can be applied to any given orbit around the BH, and to any stellar model whose oscillation eigenfrequencies $\omega_{n}$ and eigenfunctions $\xi_{n}$ are known. We consider here MESA (Paxton et al. (2011)) MAMS stars, with masses $M_{\star}\in[0.15M_{\odot},8M_{\odot}]$, and compute their oscillation properties with GYRE (Townsend & Teitler (2013)). For each star, we include modes with frequency $\omega_{n}\in[0.1\omega_{\star},20\omega_{\star}]$, where $\omega_{\star}=\sqrt{GM_{\star}/R_{\star}^{3}}$ is the typical oscillation frequency of the star. The addition of modes outside this frequency range does not significantly change the results.
@@
Figure 4: Specific energy distribution (left panel) and mass fallback rate (right panel) for MESA MAMS stars with masses $0.5M_{\odot}$, $1M_{\odot}$, $2M_{\odot}$, and $8M_{\odot}$. All cases assume $M_{\rm BH}=10^{6}M_{\odot}$ and $\beta=1.76\beta_{\rm crit}$, which equals $2.61$, $4.44$, $5.60$, $4.37$ in the four cases respectively. The energy $E$ and fallback time $T$ are normalized according to Eq. (44).
@@ 3.2 Energy distribution, fallback rate, and density profile
[…]
@@
The energy distributions are wider compared to those of the unperturbed stars. However, they do not feature a central plateau as seen in hydrodynamical simulations Lodato et al. (2009); Law-Smith et al. (2020); Ryu et al. (2020b). This is likely due the fact that we ignore the fluid self-gravity in the second stage of our model (Coughlin et al. (2016); Coughlin & Nixon (2019)). The fallback rate curves are consistent with previous results in the literature, peaking at $\mathcal{O}(10\Delta T)$ and asymptoting $\sim T^{-5/3}$ for large $T$. Performing about $10^{8}$ samples of the stellar fluid elements, the required computational time is about 1 minute in total.
In Fig. 5 we provide a visualization of the deformed stars at $t=t_{\rm TDE}$, showing their densities along an equatorial slice, for various masses and values of $\beta>\beta_{\rm crit}$. It is important to realize that the disruption time is different for each case, generally being after pericenter in small-$\beta$ cases, and before pericenter for large $\beta$. In the latter case, the stars appear less deformed, but their energy distribution is wider because $r_{p}$ is smaller. In all cases, the stars appear to be elongated not towards the BH, but instead along a direction similar to their orbital motion. This “lag” effect is due to their dynamical tidal deformation, and qualitatively matches what is observed in hydrodynamical simulations immediately before the star’s disruption (Khokhlov et al. (1993a, b)).
@@ 4 Discussion and Conclusions
The modeling of TDEs in the literature has followed two different approaches so far: on one hand, rough semi-analytical estimates; on the other hand, expensive hydrodynamical simulations. In this paper, we propose a new model that improves the accuracy of common semi-analytical arguments while still remaining computationally efficient. We achieve this by matching two different stages: stellar linear oscillation theory with free-falling particles. We determine the transition between the two stages introducing a novel TDE criterion, which is calibrated using hydrodynamical simulations. This model allows us to quickly compute the stellar energy distribution and the mass fallback rate.
We find that our deformed stellar density profiles and fallback rates are in good qualitative agreement with hydrodynamical simulations. The decomposition of the stellar deformation into eigenmodes also allows us to see which types of oscillations ($g$-modes, $f$-mode, or $p$-modes) are more excited during a TDE. We find that the fractional contribution of different types of oscillations is universal, only depending on the stellar mass, and not on the details of the orbit.

### v2

@@ Abstract
Tidal disruption events (TDEs) occur when stars pass close enough to supermassive black holes to be torn apart by tidal forces. Traditionally, these events are studied with computationally intensive hydrodynamical simulations. In this paper, we present a fast, physically motivated two-stage model for TDEs. In the first stage, we model the star’s tidal deformation using linear stellar perturbation theory, treating the star as a collection of driven harmonic oscillators. When the tidal energy exceeds a fraction $\gamma$ of the star’s gravitational binding energy (with $\gamma\sim\mathcal{O}(1)$), we transition to the second stage, where we model the disrupted material as free particles. The parameter $\gamma$ is determined with a one-time calibration to the critical impact parameter obtained in hydrodynamical simulations. This method enables fast computation of the energy distribution $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}E$ and fallback rate $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}T$, while offering physical insight into the disruption process. We apply our model to MESA-generated profiles of middle-age main-sequence stars. Our code is available on GitHub.
@@ 1 Introduction
As famously argued by Frank & Rees (1976) and Magorrian & Tremaine (1999), the TDE rate per galaxy per year is about $10^{-5}$ to $10^{-4}$, which is consistent with current observations (Donley et al. (2002); van Velzen (2018); Sunyaev et al. (2021); Yao et al. (2023)). To date, more than $100$ TDE flares have been observed in the central regions of various galaxies, and this number is expected to increase dramatically in the next few years (Bricman & Gomboc (2019)). Modern observations can now routinely capture both the rising and the decaying parts of the transient light curve (Holoien et al. (2019a); Shappee et al. (2019); van Velzen et al. (2019); Holoien et al. (2019b); Hung et al. (2020)). Once the light curve is observed, we are in principle able to infer the properties of the central BH and of the disrupted stars, e.g. BH mass and spin, stellar mass and age, impact parameter, etc. This information may be used to determine the SMBH mass function and the stellar dynamics near the galactic center (Stone & Metzger (2016); Yao et al. (2023)). In order to extract these properties from the observed data, a great amount of hydrodynamical simulations have been conducted to study the dynamics of TDE (Nolthenius & Katz (1982); Khokhlov et al. (1993a, b); Laguna et al. (1994); Diener et al. (1997); Ramirez-Ruiz & Rosswog (2009); Guillochon & Ramirez-Ruiz (2013); Law-Smith et al. (2017); Golightly et al. (2019a); Goicovic et al. (2019)). Pioneering simulations used polytropic stars models for simplicity, see e.g. Guillochon & Ramirez-Ruiz (2013). More recent simulations have instead studied TDEs of main-sequence stars and included relativistic effects, which presumably play an important role for “deep encounters” (Gafton & Rosswog (2019)), where the star travels very close to the BH horizon.
@@
The challenges in the simulation thus provide a strong motivation to come up with a simplified TDE model that is computationally cheap and capable of capturing the essential physical processes. In Lodato et al. (2009); Kesden (2012a, b); Servin & Kesden (2017); Coughlin & Nixon (2019); Golightly et al. (2019a); Coughlin & Nixon (2022), various approximation methods have been used to make predictions for TDE. For example, the star is assumed to be in hydrostatic equilibrium throughout and follow a nearly parabolic orbit until it reaches the pericenter (alternatively, the tidal radius). Then the star is disrupted instantaneously and each fluid element is modeled as a free particle, which follows its own ballistic orbit in the gravitational field of the black hole. This approximation is also known as frozen-in approximations (Lacy et al. (1982); Lodato et al. (2009)). This simplified model yields good agreement with simulations of the complete disruption of 5/3-polytropic (or equivalently, low-mass) stars. [^1] As noted in Golightly et al. (2019b), there are notable discrepancies between the numerically obtained fallback rates, and those predicted by the frozen-in approximation, for higher mass and evolved stars.
@@
where $\gamma$ is a calibration factor that can be pre-computed by fitting to the critical impact parameter obtained in hydrodynamical simulations, as it only depends on the stellar structure. The deformed density profile can be evaluated by integrating the tidal response over the star’s trajectory. When Eq. (5) is satisfied, the stellar perturbations are no longer stable, and the star is disrupted. In the second stage, we treat each fluid element inside the star as a free-falling particle, which allows us to evaluate the distribution of the orbital energy of the debris and the mass fallback rate. Even though most of the calculations in this model need to be done numerically, the whole computation only takes around one minute, much cheaper compared with the full hydrodynamical simulation.
@@ 2.2 Tidal energy and disruption criterion
The depth of the potential well is given by the gravitational binding energy of the star $U_{\rm bind}$, where
$U_{\rm bind}=-\int_{0}^{M_{\star}}\frac{Gm(r)dm}{r}~,\quad m(r)\equiv\int_{0}^{r}\rho_{0}dV~.$ (31)
We thus expect the star to be disrupted when $E_{Q}\sim|U_{\rm bind}|$. Because we extrapolate our linear oscillations model until the star is disrupted, we allow for a $\mathcal{O}(1)$ correction to the above relation. We thus introduce a calibration factor $\gamma$ and define the tidal disruption criterion as
@@ 2.3 Stage 2: free particles
$\boldsymbol{a}=-\nabla_{\boldsymbol{x}}(\Phi_{\star}+\Phi_{\rm BH})$ (37)
@@
Any function $f(E,L_{z})$ of the orbital energy $E$ and angular momentum $L_{z}$ is a solution to Eq. (36) because of the conservation laws for Newtonian orbits. In reality, it is also crucial to notice that ignoring pressure and self-gravity may lead to both qualitative and quantitative differences in the energy distribution and the fallback rate compared to numerical simulations (Fancher et al. (2023)).
In practice, we calculate the Lagrangian displacement at the disruption time $t_{\rm TDE}$ (i.e., when $E_{Q}=\gamma|U_{\rm bind}|$), which also allows us to obtain the density profile at that instant. After that, we sample the stellar fluid elements to determine the orbital energy distribution $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}E$. In Newtonian gravity, we can further use the relation
@@
to calculate the mass fallback rate $\mathop{\mathrm{d}\!}{}M/\mathop{\mathrm{d}\!}{}T$, which, to leading order, is responsible for determining the observed lightcurve in a TDE (Coughlin & Nixon (2022)).
@@ 3.1 MESA stars on parabolic orbits
The model described in Section 2 can be applied to any given orbit around the BH, and to any stellar model whose oscillation eigenfrequencies $\omega_{n}$ and eigenfunctions $\xi_{n}$ are known. We consider here MESA (Paxton et al. (2011)) MAMS stars, with masses $M_{\star}\in[0.15M_{\odot},8M_{\odot}]$, and compute their oscillation properties with GYRE (Townsend & Teitler (2013)). [^5] We define MAMS as having a hydrogen fraction of $0.35$. However, low-mass stars take much longer than the age of the Universe to become MAMS, so they are more likely to be well described as zero-age main sequence (ZAMS) stars, with a hydrogen fraction of $0.7$. For simplicity, in this paper we ignore this issue and assume all stars are MAMS. For each star, we include modes with frequency $\omega_{n}\in[0.1\omega_{\star},20\omega_{\star}]$, where $\omega_{\star}=\sqrt{GM_{\star}/R_{\star}^{3}}$ is the typical oscillation frequency of the star. The addition of modes outside this frequency range does not significantly change the results.
@@
Figure 4: Specific energy distribution (left panel) and mass fallback rate (right panel) for MESA MAMS stars with masses $0.5M_{\odot}$, $1M_{\odot}$, $2M_{\odot}$, and $8M_{\odot}$. All cases assume $M_{\rm BH}=10^{6}M_{\odot}$ and $\beta=1.76\beta_{\rm crit}$, which equals $2.61$, $4.44$, $5.60$, $4.37$ in the four cases respectively. The left panel is normalized according to Eq. (45), while the right panel displays physical units.
@@ 3.2 Energy distribution, fallback rate, and density profile
Figure 5: Equatorial density slices at $t=t_{\rm TDE}$ for different values of $M_{\star}$ and $\beta$. The top row corresponds to $\beta=1.05\beta_{\rm crit}$, the middle row to $\beta=1.76\beta_{\rm crit}$ (equivalent to the four cases described in Fig. 4), the bottom row to $\beta=5\beta_{\rm crit}$. All cases assume $M_{\rm BH}=10^{6}M_{\odot}$. In the top row, the disruption time $t_{\rm TDE}$ is after pericenter; in the bottow two rows, it is before pericenter. Note that the orientation of this figure is rotated counterclockwise by $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ with respect to Fig. 1. We denote the direction pointing to the BH with an arrow. At pericenter, the arrow points left. Each panel has a width of $4R_{\star}$ and a height of $6R_{\star}$, while the logarithmic color scale is normalized in each panel to the maximum star density.
@@
Performing about $10^{8}$ samples of the stellar fluid elements, the required computational time is about 1 minute in total.
The energy distributions are wider compared to those of the unperturbed stars. However, they do not feature a central plateau as seen in hydrodynamical simulations Lodato et al. (2009); Law-Smith et al. (2020); Ryu et al. (2020b). This is likely due the fact that we ignore the fluid self-gravity in the second stage of our model (Coughlin et al. (2016); Coughlin & Nixon (2019)). The fallback rate curves are consistent with previous results in the literature, peaking at $\mathcal{O}(10\Delta T)$ and asymptoting $\sim T^{-5/3}$ for large $T$. When comparing with the numerical simulations in (Fancher et al. (2023)), the peak timescales for the fallback rates modeled using our formalism, while being a significant improvement over the frozen-in approximation, appear to be systematically shorter. This is largely due to the abrupt transition from the perturbative stellar oscillation phase to the free-particle regime in the current model, neglecting fluid relaxation driven by pressure gradients and self-gravity.
In Fig. 5 we provide a visualization of the deformed stars at $t=t_{\rm TDE}$, for various masses and values of $\beta>\beta_{\rm crit}$. We create the figure by sampling the fluid elements on an equatorial slice of the star (uniformly in the azimuthal angle $\phi$), and computing their Lagrangian displacements as in Eq. (13) and (20). The positions of all fluid elements (weighed by their mass) are then binned to create a density map, to which we add a constant background equal to $10^{-8}\rho_{c}$ to improve visual clarity. It is important to realize that the disruption time is different for each case, generally being after pericenter in small-$\beta$ cases, and before pericenter for large $\beta$. In the latter case, the stars appear less deformed, but their energy distribution is wider because $r_{p}$ is smaller. In all cases, the stars appear to be elongated not towards the BH, but instead along a direction similar to their orbital motion. This “lag” effect is due to their dynamical tidal deformation, and qualitatively matches what is observed in hydrodynamical simulations immediately before the star’s disruption (Khokhlov et al. (1993a, b)).
@@ 4 Discussion and Conclusions
The modeling of TDEs in the literature has followed two different approaches so far: on one hand, rough semi-analytical estimates; on the other hand, expensive hydrodynamical simulations. In this paper, we propose a new model that improves the accuracy of common semi-analytical arguments while still remaining computationally efficient. We achieve this by matching two different stages: stellar linear oscillation theory with free-falling particles. We determine the transition between the two stages introducing a novel TDE criterion, which is calibrated by matching the critical impact parameter to the one observed in hydrodynamical simulations available in the literature. This model allows us to quickly compute the stellar energy distribution and the mass fallback rate.
We find that our deformed stellar density profiles and fallback rates are in good qualitative agreement with those found in the literature. The decomposition of the stellar deformation into eigenmodes also allows us to see which types of oscillations ($g$-modes, $f$-mode, or $p$-modes) are more excited during a TDE. We find that the fractional contribution of different types of oscillations is universal, only depending on the stellar mass, and not on the details of the orbit.